OpenStudy (anonymous):

Multiply these binomials. (3x - 5)(3x + 5) A. 9x2 - 15x - 25 B. 9x2 - 30x - 25 C. 6x - 25 D. 9x2 - 25

OpenStudy (anonymous):

your answer is d

OpenStudy (anonymous):

its simple (a+b)(a-b) =a^2 - b^2 so answr is option D

Parth (parthkohli):

Answer is option D, You just gotta do it -> (a+b)(a-b) = a^2 - b^2